OpenStudy (anonymous):

how could i solve this equation? 3m - 5 = 7m - 21

OpenStudy (anonymous):

7m-3m=21-5 4m=16 m=4

OpenStudy (anonymous):

what im trying to figure out is where the 4m =6 came from.

OpenStudy (anonymous):

just move 7m to left side of equal n move 5 to right side of equal.. signs ll change

OpenStudy (anonymous):

oh,i see now. thanks!

OpenStudy (anonymous):

3m - 5 = 7m - 21 3m - 7m = -21 + 5 -4m = -16 m = -16/-4 = 4

OpenStudy (anonymous):

Just isolate the variable on one side and keep changing the signs as you move the constants to the other side of the equation.